### v2

# Brainformer: Mimic Human Visual Brain Functions to Machine Vision Models via fMRI

Abstract
Human perception plays a vital role in forming beliefs and understanding reality. A deeper understanding of brain functionality will lead to the development of novel deep neural networks. In this work, we introduce a novel framework named Brainformer, a straightforward yet effective Transformer-based framework, to analyze Functional Magnetic Resonance Imaging (fMRI) patterns in the human perception system from a machine-learning perspective. Specifically, we present the Multi-scale fMRI Transformer to explore brain activity patterns through fMRI signals. This architecture includes a simple yet efficient module for high-dimensional fMRI signal encoding and incorporates a novel embedding technique called 3D Voxels Embedding. Secondly, drawing inspiration from the functionality of the brain’s Region of Interest, we introduce a novel loss function called Brain fMRI Guidance Loss. This loss function mimics brain activity patterns from these regions in the deep neural network using fMRI data. This work introduces a prospective approach to transfer knowledge from human perception to neural networks. Our experiments demonstrate that leveraging fMRI information allows the machine vision model to achieve results comparable to State-of-the-Art methods in various image recognition tasks.
1 Introduction
Recent studies in machine vision understanding [26, 28, 37, 86, 45, 15] have demonstrated the effectiveness of single-encoder models through pretraining on image datasets, e.g., ImageNet [13]. These methods are designed to acquire universal visual representations of objects that can be flexibly applied to various downstream tasks, including object detection and semantic segmentation. Nonetheless, the most significant limitation of these methods lies in the costly annotation process, mainly when applied at a large scale. In response to this challenge, self-supervised techniques have emerged [4, 25, 57]. They aim to acquire visual representations without incurring human annotation expenses while delivering commendable performance compared to supervised methods.
To avoid expensive annotation efforts, a surge in the development of foundational language models, e.g., BERT [14], GPT-2, GPT-3 [6], RoBERTa [44], T5[74], BART[39] is observed. These approaches are typically trained on extensive datasets and utilize the text as guidance for the vision models [93, 72, 29]. From the recent success of text as supervision for the visual learning, we revisit one of the key questions in the early days of artificial intelligence: What if human brain behaviors can serve as the guiding force for the machine vision models?. To answer this question, we found that Functional Magnetic Resonance Imaging (fMRI) has provided valuable insights into various aspects of human cognition and neuroscience. It contains rich information on how the human visual system works. For example, fMRI can help identify the specific regions of the brain that are active during various cognitive tasks. Studies [21, 22, 91]have shown that different intelligence-related tasks activate specific brain areas, such as problem-solving or memory. In addition, human intelligence is not just about the activity of individual brain regions but also about how different brain regions communicate. fMRI can reveal functional connectivity patterns, highlighting networks involved in various cognitive processes.
While text typically reflects the outcomes of the recognition process, fMRI captures the dynamics of the cognitive processes involved. For instance, examiners are tasked with describing the context of an image, such as “A man is sitting on the chair." Initially, they focus on the man and then shift attention to the chair, determining the interactions between these objects, specifically, the act of sitting. Subsequently, they conclude the context in the text form. When encountering this description for the first time, it remains unclear which object the examiner attended to first or whether other smaller objects influenced their perception. Notably, this information can be encoded in fMRI signals. By analyzing specific brain regions or regions of interest, we can explore the whole recognition process of the human brain. This work provides a promising approach to further studies and endeavors to bridge the gap between human intelligence and deep neural networks. The contributions of this work can be summarized as follows.
Figure 1: Given a pair of images and fMRI signals (x-axis is the voxel index, y-axis is the magnitude of the voxel response), Brainformer can explore the local patterns of fMRI signals from brain regions and discover their interactions. Best view in color
Contributions of this Work. This paper presents a new simple but efficient Transformer-based framework, named Brainformer, to analyze and leverage fMRI signals captured from human brain activities to supervise the machine vision learning model. Second, a new Brain fMRI Guidance Loss function is presented to distill information from fMRI extracted using Brainformer. It serves as a guidance signal to enhance the performance capabilities of vision models. Finally, Brainformer is designed for self-supervised learning and trained as an end-to-end deep network. It consistently outperforms previous State-of-the-Art Self-supervised learning methods across standard benchmarks, i.e., object detection, instance segmentation, semantic segmentation, and brain response prediction. The code will be released.
2 Related Work
Vision-Language Foundation Models. Pretraining on large-scale datasets [13, 50, 94] has become a popular approach for many visual recognition problems [27, 25, 59, 61, 56, 54, 55], e.g., classification, localization, or segmentation, etc. Despite this method has great success, it is limited by the number of annotated training data that is costly to collect. Recently, self-supervised training has been proposed to address this problem [60, 53, 58]. Significant advancements have occurred in vision-language pertaining [73, 29, 96, 66, 89, 87, 40, 92, 95, 49, 88]. This approach aims to encode both visual and text information into a model [82, 9, 97, 75, 36, 5, 53]. Recent work has proposed Image-Text foundation models that bridge the gap between image and text understanding [72, 30, 67, 89, 87]. They leverage dual-encoder architecture pre-trained on noise image-text pairs using contrastive learning or generative losses. Their studies showcase impressive performance in vision-language benchmarks while maintaining strong visual encoder capabilities for classification tasks.
Leveraging Human Brain Functions In Training. While the text has successfully demonstrated its efficiency in helping machine learning models enhance their performance, recent studies have been inspired by human brain mechanisms to improve models [84, 85, 32, 69, 41, 51, 18, 76, 43, 68, 34, 12, 77, 16, 38, 84, 71]. Some recent methods have used neural activity data to guide the training of models [65, 19, 80, 62, 52]. They utilized EEG and fMRI signals to constrain the neural network to behave like the neural response in the visual cortex.
Decoding functional MRI. Decoding visual information from fMRI signals has been studied for a decade [24, 83, 31, 11, 23]. Most of these studies aimed to explore the hidden information inside the brain. It is a difficult task because of the low signal-to-noise ratio. Recently, with the help of deep learning, the authors in [10, 78, 81, 64, 41, 63] proposed methods to reconstruct what humans see from fMRI signal using diffusion models. Different from recent studies on decoding fMRI [81, 78, 10], our goal is to explore valuable vision information from fMRI signal, i.e., semantic and structure information, and utilize them as supervision for helping to enhance the recognition capability of the vision model. Although our approach also aims to extract feature representations of fMRI data, these previous methods still have limitations and do not apply to our problem. First, prior approaches, such as those by Yu et al. [81] and Paul et al., [78] relied on linear models and Multi-layer Perceptron to generate fMRI feature representations. This approach cannot capture the local pattern of the signals, especially the non-linear or complicated patterns. Kim et al. [35] presented SwiFT for self-training fMRI in 4D data, which does not apply to our problem. On the other hand, MindVis [10] utilized a Masked Brain Model inspired by MAE, which explored local patterns but ignored the correlations between multiple voxels, which describes how neurons interact inside the brain. Furthermore, MindVis failed to leverage the 3D spatial information in fMRI signals, which is crucial in understanding the relationships between voxels. Most notably, none of these approaches leverage characteristics of functional regions of interest (ROI) within the brain, which contain rich information about visual stimuli.
3 The Proposed Approach
3.1 Motivation
The fMRI contains a structured and semantically rich representation of information since fMRI captures brain activities in the context of a person engaging in visual tasks. It provides ground truth data about how the human brain responds to visual stimuli. According to recent studies [70, 33, 17], the brain can be divided into several Regions of Interest (ROI), where each region holds a different function. Especially in the scope of visual cognition, these are 6 regions specified as Early retinotopic visual regions (prf-visualrois), Body-selective regions (floc-bodies), Face-selective regions (floc-faces), Place-selective regions (floc-places), Word-selective regions (floc-words) and Anatomical streams (streams). In general, these regions are the responses for processing information related to the human body, face, place, motion, or objects regarding identification, recognition, etc. Therefore, the fMRI signals from these regions can be used to train computer vision models, providing a reference for what the brain is “seeing" during the training. By training vision models jointly with brain activities observed in fMRI scans, they can explore the relationship between the visual information processed by the brain and the visual information that models are learning. It can help uncover how the brain represents and processes visual information.
3.2 Brainformer
As aforementioned, human intelligence is presented by the activity of individual brain regions of interest and how they interact with each other. Inspired by this concept, we introduce Brainformer, as shown in Figure 2, a model that takes fMRI signals from Regions of Interest in the brain as inputs. The proposed model comprises two primary modules. First, we introduce a Mutli-scale fMRI Transformer, as shown in Figure 3, designed to uncover local patterns of brain activities within each ROI. Subsequently, we feed a sequence of ROI features derived from the output of the Mutli-scale fMRI Transformer into a conventional Transformer to estimate the correlation and interaction among multiple ROIs. In the following sections, we focus on the design of the Mutli-scale fMRI Transformer, as detailed in Section 3.2.3. Section 3.2.4 introduces brain cognitive features which represent for correlation between brain regions. Prior to that, we present an efficient way to extract features of raw signals in Section 3.2.1. Additionally, we introduce a novel technique called Brain 3D Voxel Embedding in Section 3.2.2, aimed to preserve the spatial information of the signals.
Figure 2: Brainformer utilizes fMRI signals (x-axis is the voxel index, y-axis is the magnitude of the voxel response) from specific brain regions as input, extracting the local features representing patterns within each region. The TransformerBlock measures the correlation among these regions to emulate brain activities. This information is subsequently transferred to the vision model through Contrastive Loss and Brain fMRI Guidance Loss.
3.2.1 High-dimensional fMRI Feature Encoding
Figure 3: The details of Multi-scale fMRI Transformer module.
Let $m^{k}$ be the fMRI signals of the $k^{th}$ region of interest in the brain. This signal can be constructed using Eqn. (1).
$m^{k}=\left[\delta(x_{i},y_{i},z_{i})\right]_{i=0}^{N^{k}-1}$ (1)
where $\delta$ is the function that takes the value of change in blood and oxygenation in the voxel coordinated at $(x_{i},y_{i},z_{i})$. $N^{k}$ is the number of voxels in this region. This definition shows that fMRI is a 1D high-dimension signal due to the significant value of $N^{k}$, e.g., $N^{k}\approx 20K$. A straightforward approach to model this signal is adopting linear or fully connected layers [78] and extracting its latent embedding. This approach, however, has two drawbacks. First, fully connected layers with high dimensional features are inefficient as it is challenging to learn useful information from the input space while maintaining a high memory usage of the model. Second, it focuses more on the global structure while ignoring the local patterns presented in the fMRI signal. To solve these two problems, we propose Conv1D to extract features of the fMRI signal. Formally, the embedding feature of fMRI is represented in Eqn. (2).
$\begin{split}\mathbf{r}^{k}&=\texttt{Conv1D}(m^{k})\in\mathbb{R}^{N^{k}\times d_{r}}\end{split}$ (2)
The $\mathbf{r}^{k}$ is the embedding features of $m^{k}$ after the convolution step, and $d_{r}$ is the embedding dimension.
3.2.2 Brain 3D Voxel Embedding
As shown in Eqn (1), the fMRI signal is a compressed representation of Magnetic Resonance Imaging (MRI) containing detailed 3D information. However, direct learning from the flattened signal will ignore the spatial structure. Figure 4 illustrates this problem. Consider two voxels, denoted as $v_{1}=(x_{1},y_{1},z_{1})$ and $v_{2}=(x_{2},y_{2},z_{2})$, which are closely situated in the 3D space of MRI, but in fMRI signals representation, they appear distant from each other. Therefore, it becomes crucial to incorporate spatial information for the model to uncover deeper structural features. Meanwhile, prior studies [78, 81, 10] still need to address this problem. In light of this concept, we introduce a new approach named Brain 3D Voxel Embedding to capture the spatial architecture of fMRI more effectively.
Let $m_{i}^{k}$ be the voxel in the signal $m^{k}$ and $(x_{i}^{k},y_{i}^{k},z_{i}^{k})$ are the 3D coordinates of this voxel. We use Linear function to map these $(x_{i}^{k},y_{i}^{k},z_{i}^{k})$ coordinates into the same dimension, i.e., $d_{r}$ dimension, as the fMRI embedding features in Eqn. (2)
$\begin{split}v_{i}^{k}&=\texttt{Linear}(x_{i}^{k},y_{i}^{k},z_{i}^{k})\in\mathbb{R}^{d_{r}}\\ v(m^{k})&=\texttt{concat}\left[v_{0}^{k},v_{1}^{k},\dots,v_{N^{k}-1}^{k}\right]\in\mathbb{R}^{N^{k}\times d_{r}}\end{split}$ (3)
where $v_{i}^{k}$ is the brain 3D voxel embedding of single voxel $i^{th}$, $v(m^{k})$ is the embedding of the entire signal $m^{k}$. Incorporating with Brain 3D Voxel Embedding to the Eqn. (2), we get the features of fMRI as shown in Eqn. (4).
$\small\begin{split}\mathbf{r}^{k}&=\texttt{Conv1D}(m^{k})+v(m^{k})\end{split}$ (4)
3.2.3 Mutli-scale fMRI Transformer
Figure 4: Two voxels, denoted $v_{1}=(x_{1},y_{1},z_{1})$ and $v_{2}=(x_{2},y_{2},z_{2})$ that are located closely in the 3D space of MRI, but in the fMRI signals (x-axis is the voxel index, y-axis is the magnitude of the voxel response), they are far awya from each other.
While extremely powerful for a wide range of natural language processing tasks, Transformers have limitations when handling long sequences. These limitations are primarily due to the quadratic complexity of the self-attention mechanism and the model’s parameter count. In addition, the attention score might collapse and close to zero if the sequence length is considerable. Hence, using the typical transformer for lengthy fMRI signals is not a complete solution.
To deal with the long sequence in our problem, we propose the Mutli-scale fMRI Transformer, shown in Figure 3. It contains multiple levels, each level consisting of a Transformer block denoted as TransBlock. The length of the signal is reduced when passing through each level of the network until it is sufficient enough. The details of the Multi-scale fMRI Transformer are presented in Algorithm 1. In particular, each level consists of two steps:
  • Step 1: We employ a slicing window with $w$ of width that traverses the entire signal $\mathbf{r}^{k}$ from its beginning to the end, with a step size of $s$. This process decomposes $\mathbf{r}^{k}$ into smaller sub-sequences, denoted as $\mathbf{q}_{i}^{k}=\mathbf{r}^{k}\left[i*s:i*s+w\right]\quad 0\leq i\leq n_{s}=\frac{N^{k}}{s}$ that has a computation-efficiency length for the Transformer. The adjacent subsequences overlap by a distance of $s$, preventing the loss of local information, and $n_{s}$ is the number of subsequences.
  • Step 2: We fed $\mathbf{q}_{i}^{k}$ into the TransBlock to learn the patterns inside the signal and get the features $\bar{\mathbf{q}}_{i}^{k}=\texttt{TransBlock}^{k}(\mathbf{q}_{i}^{k})$. After that, all the features $\bar{\mathbf{q}}_{i}^{k}$ of subsequences are gathered to form a new sequence $\textbf{t}=\left[\bar{\mathbf{q}}_{i}^{k}\right]_{i=0}^{n_{s}-1}$ and then passed into the next level.
3.2.4 Brain Cognitive Features
Let $\left[\bar{\mathbf{q}}^{k}\right]_{k=0}^{N_{r}}$ be the list of fMRI features that are the outputs from the Mutli-scale fMRI Transformer. $N_{r}$ is the number of regions in the brain. In particular, we define $N_{r}=6$ as discussed in the section 3.1. To learn the correlation between the ROIs, we adopt another feature embedding process by feeding $\left[\bar{\mathbf{q}}^{k}\right]_{k=0}^{N_{r}}$ into a final TransBlock and receiving the brain cognitive features, denoted as $\mathbf{q}=\texttt{TransBlock}(\left[\bar{\mathbf{q}}^{k}\right]_{k=0}^{N_{r}})$
3.3 Training Objectives
Our framework is trained by a combination of Contrastive Loss and Brain fMRI Guidance Loss as $\mathcal{L}=\lambda_{con}\mathcal{L}_{con}+\lambda_{bfg}\mathcal{L}_{bfg}$ where $\lambda_{con}$ and $\lambda_{bfg}$ are the weights of the loss functions. In this paper, we select $\lambda_{con}=\lambda_{bfg}=0.5$. The details of $\mathcal{L}_{con}$ and $\mathcal{L}_{bfg}$ are described in the sections below.
3.3.1 Contrastive Loss
Let $\mathbf{p}\in\mathbb{R}^{d_{r}}$ be the image features extracted by the image encoder, e.g., SwinTransformer, ConvNext. We employ contrastive loss [8] to align visual representation with the brain cognitive features $\mathbf{q}$ as in Eqn. (5).
$\begin{split}\mathcal{L}_{con}&=-\frac{1}{N}\sum_{i}^{N}\log\frac{\exp(\mathbf{p}_{i}\otimes\mathbf{q}_{i}/\sigma)}{\sum_{j}^{N}\exp(\mathbf{p}_{i}\otimes\mathbf{q}_{j}/\sigma)}-\frac{1}{N}\sum_{i}^{N}\log\frac{\exp(\mathbf{q}_{i}\otimes\mathbf{p}_{i}/\sigma)}{\sum_{j}^{N}\exp(\mathbf{q}_{i}\otimes\mathbf{p}_{j}/\sigma)}\end{split}$ (5)
where $\sigma$ is the learnable temperature factor, $N$ is the number of samples, and $\otimes$ is the dot product.
3.3.2 Brain fMRI Guidance Loss
[…]
Algorithm 1 Mutli-scale fMRI Transformer
Besides the contrastive loss that aligns the global context of brain signals and stimuli image, the Brain fMRI Guidance Loss aims to align the local context. Since the ROI features $\bar{\mathbf{q}}^{k}$ is extracted from a particular region of the brain, e.g., floc-bodies, floc-faces, etc., it embeds the information on how the brain perceives the objects inside the image. If we can leverage these features as the guidance during training, the vision model can mimic the perception system of the human.
Let $\bar{\mathbf{p}}^{k}$ be the vision features representing a specific local context within the image, e.g., objects, persons, the background, or locations. These features can be achieved by passing pooling features of the vision model to the linear layer and then projected to the same dimension space as $\bar{\mathbf{q}}^{k}$. Our goal is to encourage the visual perceptions of the vision model to be similar to the ROI features of the brain. This requires maximizing the similarity between two features, $\bar{\mathbf{p}}^{k}$ and $\bar{\mathbf{q}}^{k}$. Furthermore, it is important to emphasize that the brain’s ROIs have distinct functions. Consequently, we must ensure the dissimilarity of $\bar{\mathbf{q}}^{k}$ with respect to all $\bar{\mathbf{q}}^{g}$, where $k\neq g$. To facilitate the two constraints, we propose the Brain fMRI Guidance Loss as in Figure 5 and in Eqn. (6).
$\begin{split}\mathcal{L}_{bfg}&=-\frac{1}{N}\sum_{i}^{N}\sum_{k}^{N_{r}}\log\frac{\exp(\bar{\mathbf{p}}^{k}_{i}\otimes\bar{\mathbf{q}}^{k}_{i})}{\sum_{g}^{N_{r}}\exp(\bar{\mathbf{p}}^{k}_{i}\otimes\bar{\mathbf{q}}^{g}_{i})}\end{split}$ (6)
4 Experimental Results
4.1 Datasets
To pretrain Brainformer, we leverage the Natural Scenes Dataset (NSD) [1], a comprehensive compilation of responses from eight subjects obtained through high-quality 7T fMRI scans. Each subject was exposed to approximately 73,000 natural scenes, forming the basis for constructing visual brain encoding models.
4.2 Brainformer Training
We use data from seven subjects in NSD for training and leave one for testing. The images are resized to 224 $\times$ 224 before feeding into the vision model. We do not use any augmentations for the image because the fMRI signals depend on the input image. Any change in vision stimuli will affect human recognition. For the fMRI signals, we divide them into six regions of interest following [20] and feed them into Brainformer simultaneously. Experimentally, we select Conv1D with kernel size and stride are 32 and 16, respectively.
Figure 5: The circle and rectangle represent vision and fMRI features, respectively. Each color indicates a different object of interest that the human brain is processing. The Brain fMRI Guidance Loss aims to align visual and fMRI features of the same object while discriminating with features of other objects.
The Multi-scale fMRI Transformer is designed with a window size as $w=64$, stride as $s=32$, and number of levels as $h=2$. Feature dimensions of image and fMRI signals are projected into $d_{r}=768$ space. For the vision model, we select Swin-S [45] and ConvNext-S [46] as the image backbone. Brainformer is easily implemented in the Pytorch framework and trained on 8 $\times$ A100 GPUs. The initial learning rate is 0.0001 and decreases gradually following the ConsineLinear policy [47]. A batch size of 64 per GPU is employed. Optimization is performed using AdamW [48] with 100 epochs, with the entire training process concluding within two hours. The pre-trained models are used for further downstream tasks, including COCO object detection [42] and ADE20K semantic segmentation [98].
Table 1: Results of object detection and instance segmentation on COCO dataset
                           (a) Object Detection
[…]
Swin-S Random init 41.3($\pm$0.2) 63.4($\pm$0.3) 45.5($\pm$0.4)
Swin-S CLIP [72] 41.9($\pm$0.1) 64.0($\pm$0.4) 46.0($\pm$0.3)
Swin-S Brainformer 43.6($\pm$0.2) 65.8($\pm$0.5) 47.4($\pm$0.3)
ConvNext-S Random init 42.6($\pm$0.3) 65.5($\pm$0.2) 47.0($\pm$0.2)
ConvNext-S CLIP [72] 42.8($\pm$0.1) 66.1($\pm$0.2) 48.3($\pm$0.3)
ConvNext-S Brainformer 45.1($\pm$0.1) 68.2($\pm$0.2) 50.0($\pm$0.2)
(b) Semantic Segmentation
[…]
Swin-S Random init 38.4($\pm$0.2) 60.8($\pm$0.3) 41.3($\pm$0.2)
Swin-S CLIP[72] 39.2($\pm$0.3) 61.3($\pm$0.3) 41.7($\pm$0.2)
Swin-S Brainformer 43.1($\pm$0.2) 63.1($\pm$0.5) 43.3($\pm$0.2)
ConvNext-S Random init 39.8($\pm$0.4) 61.7($\pm$0.1) 42.9($\pm$0.3)
ConvNext-S CLIP [72] 40.8($\pm$0.4) 62.2($\pm$0.2) 43.5($\pm$0.1)
ConvNext-S Brainformer 44.0($\pm$0.2) 64.0($\pm$0.5) 45.3($\pm$0.4)
4.3 Object Detection on COCO
Settings. We conducted object detection and instance segmentation experiments using the COCO 2017 dataset, which comprises 118,000 training images, 5,000 validation images, and 20,000 test-dev images. Notably, we excluded images from the NSD dataset to prevent any data leakage, a subset of COCO. For these experiments, we utilized the MaskRCNN framework [3] implemented with mmdetect [7] for enhanced performance efficiency.
Performance. The object detection and instance segmentation results are presented in Table 1. In summary, our approach, which leverages fMRI signals for training, consistently outperforms the CLIP framework, where models rely solely on textual supervision. Specifically, when comparing Swin-S/CLIP to Swin-S/Random, the Swin-S/CLIP achieves approximately a 0.6% improvement in box/AP and a 0.8% improvement in seg/AP. However, our method Swin-S/Brainformer presents even better performance, surpassing Swin-S/CLIP by approximately 1.7% and 3.9% in object detection and instance segmentation, respectively, translating to a substantial 2.3% and 4.7% improvement compared to the same model Swin-S/Random without any pretraining method. We also observe the same results with the ConvNext-S backbone, where Brainformer achieves around 2.3% and 3.2% higher than CLIP in box/AP and seg/AP.
4.4 Semantic Segmentation on ADE20K
Settings. For semantic segmentation, we conducted training using UpperNet [90] on the ADE20K database, which includes a wide spectrum of 150 semantic categories. This dataset comprises 25,000 images, with 20,000 allocated for training, 2,000 for validation, and 3,000 for testing.
Performance. Table 2 presents the semantic segmentation results. From these results, Swin-S/Brainformer is $1.48\%$ mIoU (41.77 v.s 40.29) higher than Swin-S/CLIP while ConvNext-S/Brainformer maintains the performance better than ConvNext-S/CLIP by $1.65\%$ mIoU.
Table 2: Results of semantic segmentation on ADE20K and brain activities response prediction on NSD.
  ADE20K NSD
Backbone Pretrain mIoU PCC
Swin-S Random init 38.37($\pm$0.2) 40.41($\pm$0.4)
Swin-S CLIP [72] 40.29($\pm$0.3) 41.25($\pm$0.3)
Swin-S Brainformer 41.77($\pm$0.3) 44.63($\pm$0.2)
ConvNext-S Random init 39.22($\pm$0.3) 54.21($\pm$0.3)
ConvNext-S CLIP [72] 41.27($\pm$0.4) 55.70($\pm$0.2)
ConvNext-S Brainformer 42.92($\pm$0.3) 57.43($\pm$0.4)
4.5 Human Brain Response Prediction NSD
Settings. This experiment aims to predict the human brain response to complex natural scenes, as recorded during participants’ observations [20, 2]. The dataset contains responses from eight subjects in the NSD. We use seven subjects for pretraining CLIP and Brainformer, while one is reserved for the downstream task. We follow the evaluation protocols outlined in [20], utilizing the Pearson Correlation Coefficient (PCC) score as our evaluation metric.
Performance. The results of our brain response prediction are shown in Table 2. Significantly, Brainformer yields substantial performance improvements in this task. Specifically, Swin-S/Brainformer demonstrates approximately 4.22% and 3.38% better performance than Swin-S/Random and Swin-S/CLIP, respectively. Furthermore, it is worth noting that ConvNext-based models perform strong predictive capabilities, with ConvNext-S/Brainformer achieving a PCC of 57.43%, approximately 1.73% higher than ConvNext-S/CLIP.
5 Ablation Studies
Table 3: Performance on various settings.
  COCO ADE20K NSD
[…]
pos embed 41.9 41.8 41.72 56.19
3D voxel embed 45.1 44.0 42.92 57.43
w/o Brain fMRI Guidance Loss 42.7 40.7 41.80 56.25
w/ Brain fMRI Guidance Loss 45.1 44.0 42.92 57.43
$w=128,s=64$ 41.9 41.1 40.90 55.80
$w=128,s=32$ 42.3 41.6 41.80 56.20
$w=64,s=32$ 45.1 44.0 42.92 57.43
##\## subjects = 1 42.5 39.3 39.15 54.20
##\## subjects = 3 42.9 39.6 39.37 54.51
##\## subjects = 5 43.4 40.4 40.08 55.01
##\## subjects = 7 45.1 44.0 42.92 57.43
In this section, we study the efficiency of the Brain 3D Voxel Embedding as presented in Section 3.2.2, Brain fMRI Guidance Loss in Section 3.3.2, hyper-parameters of Multi-scale fMRI Transformer in Section 3.2.3, and performance of Brainformer on different amounts of data. We select ConvNext-S as the backbone for these ablation studies.
Brain 3D Voxel Embedding. In Table 3, we provide an ablation study for 3D voxel embedding. The use of this embedding yields an improvement of approximately +3.2% box/AP, +2.2% seg/AP, +1.2% mIoI, and +1.24%PCC over the network employing traditional positional embedding in various tasks such as object detection, instance segmentation, semantic segmentation, and brain response prediction, respectively. These findings demonstrate the effectiveness of the 3D voxel embedding for fMRI signals.
Brain fMRI Guidance Loss. We also provide an ablation study for the usage of Brain fMRI Guidance Loss in Table 3. The model trained with the guidance loss performs better than not using it. In particular, this loss function helps to improve +2.4% box/AP, +3.3% seg/AP, +1.12% mIoU, and 1.18%PCC for object detection, instance segmentation, semantic segmentation, and brain response prediction, respectively. We also investigate how Brain fMRI Guidance Loss facilitates the transfer of semantic and structured information from fMRI signals to the vision model. To accomplish this, we follow a three-step process. First, we extract the fMRI features and image features from Brainformer and the vision model, respectively. Second, we compute the cosine similarity between these feature vectors. Finally, we leverage GradCam, as described in [79], by analyzing the backward gradient of the similarity to generate an attention map. This map highlights specific image regions strongly correlated to the fMRI signals. We visualize this attention map in Figure 6. It illustrates the effectiveness of the Brain fMRI Guidance Loss in transferring brain activities from the fMRI signals to the vision model.
Figure 6: Visual attention with respect to fMRI signals. The first row is the input images. The second row is the corresponding attention maps of the vision model training with CLIP. The third row is the results of training with Brainformer without Brain fMRI Guidance Loss. The last row is the results with Brainformer and guidance of Brain fMRI Guidance Loss. The warmer colors, the higher attention. Best view in color.
Hyper-parameters for Multi-scale fMRI Transformer We studied the effectiveness of the hyperparameters of Brainformer, i.e., window size and stride, on the overall performance. As shown in Table 3, the results highlight that smaller window sizes and strides lead to better performance. Increasing the window size from 64 to 128 resulted in a 2.8% decrease in box/AP, a 2.4% decrease in seg/AP, a 1.12% decrease in mIoU, and a 1.23% decrease in PCC for object detection, instance segmentation, semantic segmentation, and brain response prediction, respectively. This is attributed to the fact that a smaller window size allows Brainformer to capture more local information. Similarly, we observed similar results when keeping the window size constant and increasing the stride from 32 to 64, with larger strides causing the model to miss more information potentially.
Performance on different amounts of data. We investigate the performance of Brainformer with respect to the amount of training data. We train the Brainformer with a different number of subjects included in the pre-training step. The performance is reported in Table 3. It is clear that when we use only one subject data, the performance is not much improved compared to the random initialization. However, when we increased the number of subjects, the performance of Brainformer also increased accordingly. With respect to 7 subjects’ training data, the performance is boosted by +2.6% for box/AP, +4.7% for seg/AP, +3.77% for mIoU, and 3.23% for PCC.
6 Conclusions and Discussions
In this paper, we have investigated the feasibility of transferring human brain activities to the vision models via fMRI signals. In our proposed Brainformer, we introduce the concept of the fMRI feature technique that explores the local patterns of the signals. Additionally, the Brain 3D Voxel Embedding is proposed to preserve the 3D information that fMRI signals have been missed. The Multi-scale fMRI Transformer is presented to learn the features of a long signal. We also introduce Brain fMRI Guidance Loss inspired by the brain’s mechanisms to transfer the vision semantic information of humans into the vision models. The empirical experiments on various benchmarks demonstrated that Brainformer is competitive with another SOTA method that uses text modality for knowledge transfer.
Limitations and Future Works. Inspired by prior studies in neuroscience, we have developed our approach and demonstrated the efficiency of fMRI for training vision models. However, the proposed approach can potentially consist of minor limitations. First, although the NSD dataset is one of the large-scale datasets in the neuroscience field, compared to large-scale datasets such as ImageNet [13] in the computer vision field, this data is relatively small due to time and human efforts. However, we have demonstrated the performance of Brainformer with respect to the number of training data as in Section 6. The result shows that if the amount of data is sufficient, the performance of Brainformer can be better and potentially surpass the previous methods. Second, the primary goal of this study is to present a new perspective on how to involve human recognition mechanisms in training vision learning models. The experimental configurations such as utilizing CLIP [72], ConvNext [46], SwinTransformer [45] or even methods for downstream tasks, i.e., object detection, instance segmentation, semantic segmentation, and human brain response prediction, are conducted to prove our hypothesis on a fair basis. We leave further experiments with other settings for future studies.
Broader Impacts. Brainformer has a bi-directional impact on computer vision and the neuroscience field. In particular, we have illustrated how brain activities could help the vision model. Besides, Brainformer has potential benefits for neuroscientists to study brain activities, especially human cognition. Given a pair of images and corresponding fMRI, neuroscientists can utilize Brainformer as a valuable tool to explore which neurons inside the brain are highly activated with respect to the input image or particular objects inside, thus uncovering potential novel patterns of the brain.